OpenStudy (anonymous):

A sports company conducted a road test of a new model of a geared bike. The test rider cycled (3x - 2) miles on a flat road, (x2 - 5) miles uphill, and (2x + 7) miles downhill. Which simplified expression is equivalent to the total distance, in miles, for which the bike was tested? x2 - x x2 + 5x x2 − x + 14 x2 + 5x + 14

OpenStudy (anonymous):

(5x2 + 3x + 4) − (2x2 + 5x − 1) = 5x2 + 3x + 4 − 2x2 - 5x + 1 = 5x2 - 2x2 + 3x - 5x + 4 + 1 = 3x2 - 2x + 5. A = 3, B = -2, C = 5 2) Total Distance = Flat Road Distance + Uphill Distance + Downhill Distance = (3x - 2) + (x2 - 5) + (2x + 7) = x2 + 3x + 2x - 2 - 5 + 7 = x2 + 5x